Mauro Products distributes a single product, a woven basket whose selling price is $17 and whose variable expense is $11.9 per unit. The company’s monthly fixed expense is $8,670. Required:

1. Solve for the company’s break-even point in unit sales using the equation method. (Do not round your intermediate calculations.)

2. Solve for the company’s break-even point in dollar sales using the equation method and the CM ratio. (Do not round intermediate calculations. Round "CM ratio percent" to nearest whole percent.)

3. Solve for the company’s break-even point in unit sales using the formula method. (Do not round your intermediate calculations.)

4. Solve for the company’s break-even point in dollar sales using the formula method and the CM ratio. (Do not round intermediate calculations. Round "CM ratio percent" to nearest whole percent.)
Answer #1

1. At breakeven, Fixed Cost = Unit Sales x (Sales Price - Variable Cost)
$8670 = Unit Sales x ($17-11.90)
Unit Sales = 1700 units

2. At Breakeven, Sales = Break even units x Selling Price
= 1700 x $17 = $28900
CM Ratio = (17-11.9)/17 = 30%

3. Break Even Point (Units) = Fixed Costs / Contribution margin per unit
= $8670 / $5.10 = 1700 units

4. At Breakeven Dollar Sales = Break even units x Selling Price
= 1700 x $17 = $28900
CM Ratio = (17-11.9)/17 = 30%